OpenStudy (anonymous):

An equation is shown below: -2(4x - 1) - 7 = 5 Which statement shows a correct next step in solving the equation? The equation can become -2(4x - 1) = -2 by applying the distributive property. The equation can become -2(4x - 1) = 12 by applying the addition property of equality. The equation can become -2(4x - 1) = 12 by applying the commutative property of addition The equation can become -2(4x - 1) = -2 by applying the subtraction property of equality

OpenStudy (candy13106):

the first one

OpenStudy (anonymous):

Not the first one.

OpenStudy (anonymous):

Should be the second or third. You have to add 7 to 5, so you have 12.

OpenStudy (anonymous):

why

OpenStudy (candy13106):

how come

OpenStudy (anonymous):

I agree with wrstlr. I would go for the 3rd option.

OpenStudy (candy13106):

but u have to distrubite first

OpenStudy (anonymous):

Options 1 and 4 are frankly, impossible, since once take -7 onto the other side, it becomes +7. +7+5 = 12.

OpenStudy (anonymous):

you*

OpenStudy (anonymous):

wich one is

OpenStudy (candy13106):

idk im confused still

OpenStudy (anonymous):

The first step is to add 7 to each side. so -2(4x - 1) - 7 = 5 becomes (-2(4x - 1) - 7)+7 = (5)+7. That gets rid of the 7 on the left side and 5 on the right becomes 12. So, -2(4x - 1) = 12 The next step would be to multiply the -2 through